40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 23 17:56:13 2014, Jenne, EricQ, Update, General, Vent prep for SRC length change 
    Reply  Wed Sep 24 08:39:32 2014, Jenne, EricQ, Update, General, Vent prep for SRC length change 
       Reply  Wed Sep 24 08:40:29 2014, Steve, Update, VAC, Vent has started vent77.pngbvent77.png
Message ID: 10529     Entry time: Wed Sep 24 08:39:32 2014     In reply to: 10528     Reply to this: 10530
Author: Jenne, EricQ 
Type: Update 
Category: General 
Subject: Vent prep for SRC length change 

Quote:

As Q mentioned in elog 10527, (prompted by Koji's email this afternoon) we are prepping the IFO for vent.  Here is a copy of the pre-vent checklist from the wiki, updated as we work:

 

Pre-vent checklists

 
  1. Center all oplevs/IPPOS/IPANG
  2. Align the arm cavities for IR and align the green lasers to the arms.
  3. Make a record of the MC pointing
  4. Align the beam at the PSL angle and position QPDs
  5. Reduce input power by adjusting wave plate+PBS setup on the PSL table BEFORE the PMC. (Using the WP + PBS that already exist after the laser.)
  6. Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
  7. Close shutter of PSL-IR and green shutters at the ends
  8. Make sure the jam nuts are protecting bellows

Notes:

1 & 2:  Locked arms on IR, ran ASS.  Unlocked IFO, aligned PRM for good POP flashes, aligned SRM for symmetric AS flashes.  Aligned all oplevs.  Used PZTs to align Xgreen to arm. Used knobs to align Ygreen to arm.  With PS:L green shutter closed, Xgreen  = 0.520, Ygreen = 0.680.

3:  Moved MC servo output cable that goes to ADC from OUT2 (which we had been using for monitoring AO path signals) back to its usual OUT1 (which is MC_L).  This is used in the spot position measurement script.  Spots at:  [2.32, -0.50, 1.97, -1.11, 0.26, -1.86] mm.

4: Done -Q

5:  Removed a PD that was monitoring the light coming backwards through the Faraday that sits just after the laser, just in case (confirmed that beam dump behind PD was catching beam).  Other port of PBS just had regular black hole dump.  Adjusted half wave plate until we had ~90mW just before injection into the vacuum.

6: Completed. Locked MC manually at transmission of ~1150, but low power autolocker isn't working. This isn't a critical thing, and can be fixed at any point during the vent. -Q

7: Shutters closed. Ready for Steve to check nuts and begin venting! -Q

 

ELOG V3.1.3-